1.The mean selling price of new homes in a small town over a year was $115,000. The population standard deviation was $25,000. A random sample of 100 new home sales from this city was taken.


a. What is the probability that the sample mean selling price was more than $112,000? (1)


b. What is the probability that the sample mean selling price was between $110,000 and $120,000? (1)


c. What is the probability that the sample mean selling price was between $114,500 and $115,500? (1)


2. According to a survey, only 18% of customers who visited the web site of a major retail store made a purchase. Random samples of size 45 are selected.


a. What is the mean of all the sample proportions of customers who will make a purchase after visiting the web site? (1)


b. What is the standard deviation of all the sample proportions of customers who will make a purchase after visiting the web site?


c. What proportion of all possible samples will have between 20% and 30% of customers who will make a purchase after visiting the web site? (1)


d. What proportion of all possible samples will have less than 15% of customers who will make a purchase after visiting the web site? (1)


e. 90% of the samples will have less than what percentage of customers who will make a purchase after visiting the web site? (1)


3. Assume that house prices in a neighborhood are normally distributed with a standard deviation of $20,000. A random sample of 16 observations is taken. What is the probability that the sample mean differs from the population mean by more than $4,000? (1)


4. A manufacturer of power tools claims that the mean amount of time required to assemble their top-of-the-line table saw is 80 minutes with a standard deviation of 20 minutes. Suppose a random sample of 64 purchasers of this table saw is taken. What is the probability that the sample mean will be between 78 and 85 minutes? (1)


1
Expert's answer
2021-03-12T07:43:29-0500

1. Let "X=" the sample mean selling price: "X\sim N(\mu, \sigma^2/n)"

Given "\mu=115000, \sigma=25000, n=100."

a.

"P(X>112000)=1-P(X\leq 112000)"

"=1-P(Z\leq\dfrac{112000-115000}{25000/\sqrt{100}})"

"=1-P(Z\leq -1.2)\approx0.88493"

b.

"P(110000<X<120000)"

"=P(X<120000)-P(X\leq 110000)"

"=P(Z<\dfrac{120000-115000}{25000/\sqrt{100}})"

"-P(Z\leq\dfrac{110000-115000}{25000/\sqrt{100}})"

"=P(Z<2)-P(Z\leq -2)"

"\approx0.97725-0.02275\approx0.95450"

c.

"P(114500<X<115500)"

"=P(X<115500)-P(X\leq 114500)"

"=P(Z<\dfrac{115500-115000}{25000/\sqrt{100}})"

"-P(Z\leq\dfrac{114500-115000}{25000/\sqrt{100}})"

"=P(Z<0.2)-P(Z\leq -0.2)"

"\approx0.57926-0.42074\approx0.15852"



2. Given "n=45, p=0.18"

a. "mean=p=0.18"


b. "\sigma=\sqrt{\dfrac{p(1-p)}{n}}=\sqrt{\dfrac{0.18(1-0.18)}{45}}=0.05727"


c. The sampling distribution of the sample statistics can be considered approximately normal.


"P(0.2<X<0.3)=P(X<0.3)-P(X\leq 0.2)"

"=P(Z<\dfrac{0.3-0.18}{0.05727})-P(Z<\dfrac{0.2-0.18}{0.05727})"

"\approx P(Z<2.09534)-P(Z\leq0.34922)"

"\approx0.98193-0.63654\approx0.3454"



d.

"P(X<0.15)=P(Z<\dfrac{0.15-0.18}{0.05727})"

"\approx P(Z<-0.52383)\approx0.3002"

e.

"P(X<x)=P(Z<\dfrac{x-0.18}{0.05727})=0.9"

"\dfrac{x-0.18}{0.05727}\approx1.28155"

"x\approx0.2534"

25.34%


3.

"z=\dfrac{4000}{20000/\sqrt{16}}=0.8"

"P(Z<-0.8)=0.211855"

"P(Z<0.8)=0.788145"

"P(Z<-0.8\ or \ Z>0.8)=0.211855+(1-0.788145)"

"=0.211855+(1-0.788145)=0.42371"

4.

"P(78<X<85)"

"=P(X<85)-P(X\leq 78)"

"=P(Z<\dfrac{85-80}{20/\sqrt{64}})-P(Z\leq\dfrac{78-80}{20/\sqrt{64}})"

"=P(Z<2)-P(Z\leq -0.8)"

"\approx0.97725-0.21186\approx0.7654"